OpenStudy (anonymous):

Find the zeros of (-3(x^2)(e^-2x))+(3x(e^-2x))

OpenStudy (anonymous):

\[-3x^2e^{-2x}+3xe^{-2x}=0\\ -3xe^{-2x}\left(x-1\right)=0\] \(e^{-2x}\) is always positive, so you can divide both sides by it: \[-3x(x-1)=0\]